# Jiahao Huo1,3, Yibo Yan1,2, Boren Hu1,2, Yutao Yue1,2, Xuming Hu1,2

### v1

@@ Abstract
Projecting visual features into word embedding space has become a significant fusion strategy adopted by Multimodal Large Language Models (MLLMs). However, its internal mechanisms have yet to be explored. Inspired by multilingual research, we identify domain-specific neurons in multimodal large language models. Specifically, we investigate the distribution of domain-specific neurons and the mechanism of how MLLMs process features from diverse domains. Furthermore, we propose a three-stage framework for language model modules in MLLMs when handling projected image features, and verify this hypothesis using logit lens. Extensive experiments indicate that while current MLLMs exhibit Visual Question Answering (VQA) capability, they may not fully utilize domain-specific information. Manipulating domain-specific neurons properly will result in a 10% change of accuracy at most, shedding light on the development of cross-domain, all-encompassing MLLMs in the future. Our code will be released upon paper notification.
@@
Jiahao Huo1,3, Yibo Yan1,2, Boren Hu1,2, Yutao Yue1,2, Xuming Hu1,2thanks:  Corresponding Author 1 The Hong Kong University of Science and Technology (Guangzhou) 2 The Hong Kong University of Science and Technology, 3 Tongji University {jiahaohuotj, yanyibo70, huboren99}@gmail.com, {yutaoyue, xuminghu97}@hkust-gz.edu.cn
@@ 1 Introduction
Specifically, we investigate the activation patterns of neurons in MLLMs’ feed-forward network (FFN) layers across corpora from five distinct domains, identifying less than 1% as domain-specific neurons. The datasets we utilized include LingoQA Marcu et al. (2023), RS-VQA (HR) Lobry et al. (2020), PMC-VQA Zhang et al. (2023b), DocVQA Mathew et al. (2021) and VQAv2 Goyal et al. (2017), covering domains such as Auto Driving, Remote Sensing, Medicine, Document, and Common Scenes. Figure 2 highlights the clustering and separation of image features across the domains. Image examples of these domains can also be found in Appendix A. Based on our experiment results, we argue that differences exist among these visual domains and that the vision encoder and LLM modules in MLLMs exhibit distinct patterns for these domains. Furthermore, we propose a three-stage framework based on the distribution of domain-specific neurons among MLLM’s LLM layers, where post-projection visual features are processed by LLM. To validate our hypothesis, we employ logit lens nostalgebraist (2020) to decode the hidden states of LLM’s intermediate layers to visualize the feature transformation within transformer models Vaswani et al. (2017).
@@ 2.1 Neuron Analysis
Neuron analysis has been recently widely explored in computer vision and natural language processing, which views neuron activation as the recall of learned knowledge Mu and Andreas (2020); Sajjad et al. (2022)Bau et al. (2017) propose to automatically inspect the functionality of each visual neuron in CNNs by evaluating the alignment between individual hidden units. Hernandez et al. (2021); Oikarinen and Weng (2022); Bai et al. (2024) further extend this method to open-ended by labeling hidden neurons in visual models with natural language descriptions. Neuron analysis has also been adopted to analyze language models, including the ability of sentiment analysis Radford et al. (2017), machine translation Mu et al. (2024), knowledge storing Dai et al. (2022); Zhao et al. (2024b); Chen et al. (2024) and task solving Wang et al. (2022). Recent research has associated specific neurons in LLMs with their multilingual ability, describing these neurons as language-specific neurons Tang et al. (2024); Zhao et al. (2024c). Inspired by their work, we further expand this idea to the multimodal domain, being the first to analyze the domain-specific neurons in MLLMs.
@@ 3.2 Domain-Specific Neuron Selection
Our selection method is based on Tang et al. (2024). For each domain $D_{i},i=1,2,...,k$, we feed its image-text corpus into MLLM, and record the activated frequency of each neuron $u$ as well as the total token nums $N_{u,i}$ [^1] Note that neurons in the vision encoder and language model may receive different numbers of tokens, since projected image features are concatenated with language embeddings before being fed into language model.. The activation probability of a neuron $u$ in domain $D_{i}$ is denoted as:
$p_{u,i}=\frac{M_{u,i}}{N_{u,i}}.$ (4)
We denote the probability distribution of neuron $u$ across all domains as $P_{u}$:
@@ 3.3 Latent Embeddings Interpretation
As shown in Figure 4, the logit lens decodes the hidden states of the transformer’s intermediate layers into the distribution over the vocabulary, which can be used to interpret the model’s latent embeddings nostalgebraist (2020). Ideally, the decoded distribution converges monotonically toward the next token predicted by the model.
@@ 4.1.2 Dataset and Metrics
We select five datasets representing five different domains, namely, VQAv2 Goyal et al. (2017) for common scenes, PMC-VQA Zhang et al. (2023b) for Medical domain, DocVQA Mathew et al. (2021) for Document domain, LingoQA Marcu et al. (2023) for Auto Driving domain and RS-VQA Lobry et al. (2020) for Remote Sensing domain. For LingoQA, visual instruction for each question includes multiple images, as shown in Figure 13(a). We prepare image-question pairs of nearly the same token numbers for each domain during identifying, around 20 million tokens in LLaVA-NeXT. During evaluation, the scale of the validation set is aligned with LingoQA to make a fair comparison. For DocVQA, we report Average Normalized Levenshtein Similarity (ANLS) score Biten et al. (2019) followed by the official benchmark. For LingoQA, we use the score of Lingo-Judge Marcu et al. (2023) with the official implementation. For all other datasets, we report the top-1 accuracy (%) as the metric.
@@ 4.1.3 Implementation Details
We adhere to the default prompt templates from the official repository or the original paper during evaluation, with an additional role description for the auto-driving scenes. For more details, please refer to Appendix B. We perform the forward pass without padding or truncation during the identification process. When evaluating models across different datasets, we employ beam search with max_length of 512 and num_beams of 5 to generate answers. The temperature and length_penalty arguments are set as 0.9 and -1, respectively.
@@ 4.2.1 Distribution of Domain-specific Neurons
We identify domain-specific neurons using the method described in Section 3.2. Since neurons in different modules may have different activation patterns, as shown in Appendix C, we detected those domain-specific neurons module by module. Figure 5 shows the distribution of domain-specific neurons for each layer in each module of MLLMs.
Three-stage framework of LLM understanding multimodal features.
Two obvious turning points can be observed in both LLaVA-NeXT and InstructBLIP’s language model, one in the intermediate layer and the other near the output layer. Inspired by Zhao et al. (2024c), we thus propose a three-stage framework of LLM understanding multimodal features: 1) In the first several layers, projected features are further aligned with word space. Around the turning point, the multimodal features are embedded into a uniform representation space, where included domain-specific information needs to be processed by more domain-specific neurons. 2) Transitioning into the second phase, features are further generalized and understood by language models, where domain-specific neurons decrease sharply. 3)In the third stage, language models generate responses to the input, showing a rise of neurons specific to target tasks.
Our hypothesis aligns with the previous conclusion on smaller multimodal models like GPT-J Wang and Komatsuzaki (2021), as Schwettmann et al. (2023) argue that outputs of the projection layer are further translated in deeper layers within the transformer. To further validate our hypothesis, we employ logit lens to visualize the transformation of multimodal features within language models in Section 4.2.3.
Domain-specific information in different semantic levels.
Domain-specific neurons are mainly distributed in shallow and intermediate layers within MLLMs’ vision encoders. Prior research discussed the correlation between the semantic level and layer depth, which found that more deep layers will focus on higher-level concepts in visual networks. In our settings, the document domain contains more low-level concepts, such as line and shape, while the remote sensing and medical domain may include more high-level concepts, like architectures and organs. Therefore, document neurons are mainly gathered in bottom layers close to the input end. Another interesting phenomenon is the rise of auto driving neurons near the output layer of InstructBLIP’s Q-Former, we conjecture this may reflect the struggle of model to understand the language instructions of auto driving domain.
Table 1 demonstrates the number of neurons in each domain. Remote sensing neurons have the largest proportion in LLaVA-NeXT’s vision encoder, MLP projector and language model, while in InstructBLIP, the domain owns most specific neurons are document, auto driving and auto driving separately. We argue that the number of specific neurons reflects the understanding ability of MLLM in the target domain, as more specific neurons may mean more demanding to process domain-specific information. In contrast, less specific neurons mean more generalized features in the target domain Tang et al. (2024). In this way, we find that there exists a large visual gap between domains like remote sensing, document and medical, comparing the two domains left. Moreover, InstructBLIP seems less proficient in processing questions from auto driving, as neurons of this domain exhibit the highest number in Q-Former and LLM. There is also a similar pattern in its language model as for the auto driving domain. In other words, while visual features of auto driving domain can be processed well by existing vision encoder, the language instruction of this domain may be hard to handle for language model.
@@ 4.2.3 Case Study
To investigate how MLLM’s language model processes image tokens, we employ logit lens nostalgebraist (2020) to decode the hidden states of the language model’s intermediate layers into the probability of the next token across the vocabulary. As displayed in Figure 6, when feeding a remote sensing image-question pair into InstructBLIP, we get that the most likely token next to the second image token is ’</s>’, while the most likely token next of the last text token is the correct answer, ’no’. Interestingly, two place names, "Hermann" and "Baltimore", have appeared among the top token candidates when the image input is a remote sensing picture of New York. In multilingual literature, similar phenomena have also been observed. For instance, when Llama 2 receives the French token ’fleur’ in the input, the English concept ’__flower’ will appear in the intermediate distribution Wendler et al. (2024). This suggests that the decoded vocabulary distribution can to some extent reflect the semantic concepts understood by the language model. Despite this observation, we note that the decoded distribution of image tokens is far more sparse than text tokens; even in the output layer, the probability of the most likely next token ’</s>’ is lower than 40%. It indicates that projected tokens may be treated as a sparse mixture of concepts in the representation space instead of a simple word. We also demonstrate more cases of logit lens in different domains in Appendix D.
@@ 6 Limitations
Appendix A Visual Domain Definition
@@
Appendix B Prompt Template
B.1 Instructions templates for VQA
@@
For instructions with options, we separate options in alphabetical order, as shown in Appendix B.2. $\star$ : A role description has been provided to help models better understand the tasks in auto driving. As shown below:
@@ B.2 Prompt Examples
We display the prompt format we use for evaluation in LLaVA-NeXT, as shown in Figure 8. The prompt for InstructBLIP come from direct format in Table B.1.
Appendix C Silent Neurons in MLLM’s Vision Encoder
@@
Appendix D Logit Lens Cases

### v2

@@ Abstract
Projecting visual features into word embedding space has become a significant fusion strategy adopted by Multimodal Large Language Models (MLLMs). However, its internal mechanisms have yet to be explored. Inspired by multilingual research, we identify domain-specific neurons in multimodal large language models. Specifically, we investigate the distribution of domain-specific neurons and the mechanism of how MLLMs process features from diverse domains. Furthermore, we propose a three-stage mechanism for language model modules in MLLMs when handling projected image features, and verify this hypothesis using logit lens. Extensive experiments indicate that while current MLLMs exhibit Visual Question Answering (VQA) capability, they may not fully utilize domain-specific information. Manipulating domain-specific neurons properly will result in a 10% change of accuracy at most, shedding light on the development of cross-domain, all-encompassing MLLMs in the future. The source code is available at this URL.
@@
Jiahao Huo1,3, Yibo Yan1,2, Boren Hu1,2, Yutao Yue1,2, Xuming Hu1,2thanks:  Corresponding Author 1 The Hong Kong University of Science and Technology (Guangzhou) 2 The Hong Kong University of Science and Technology, 3 Tongji University {jiahaohuotj, yanyibo70, huboren99}@gmail.com, {yutaoyue, xuminghu}@hkust-gz.edu.cn
@@ 1 Introduction
Specifically, we investigate the activation patterns of neurons in MLLMs’ feed-forward network (FFN) layers across corpora from five distinct domains, identifying less than 1% as domain-specific neurons. The datasets we utilized include LingoQA Marcu et al. (2023), RS-VQA (HR) Lobry et al. (2020), PMC-VQA Zhang et al. (2023b), DocVQA Mathew et al. (2021) and VQAv2 Goyal et al. (2017), covering domains such as Auto Driving, Remote Sensing, Medicine, Document, and Common Scenes. Figure 2 highlights the clustering and separation of image features across the domains. Image examples of these domains can also be found in Appendix B. Based on our experiment results, we argue that differences exist among these visual domains and that the vision encoder and LLM modules in MLLMs exhibit distinct patterns for these domains. Furthermore, we propose a three-stage framework based on the distribution of domain-specific neurons among MLLM’s LLM layers, where post-projection visual features are processed by LLM. To validate our hypothesis, we employ logit lens nostalgebraist (2020) to decode the hidden states of LLM’s intermediate layers to visualize the feature transformation within transformer models Vaswani et al. (2017).
@@ 2.1 Neuron Analysis
Neuron analysis has been recently widely explored in computer vision and natural language processing, which views neuron activation as the recall of learned knowledge Mu and Andreas (2020); Sajjad et al. (2022)Bau et al. (2017) propose to automatically inspect the functionality of each visual neuron in CNNs by evaluating the alignment between individual hidden units. Hernandez et al. (2021); Oikarinen and Weng (2022); Bai et al. (2024) further extend this method to open-ended by labeling hidden neurons in visual models with natural language descriptions. Neuron analysis has also been adopted to analyze language models, including the ability of sentiment analysis Radford et al. (2017), machine translation Mu et al. (2024), knowledge storing Dai et al. (2022); Zhao et al. (2024b); Chen et al. (2024) and task solving Wang et al. (2022). Recent research has associated specific neurons in LLMs with their multilingual ability, describing these neurons as language-specific neurons Tang et al. (2024); Zhao et al. (2024c). Inspired by their work, we further expand this idea to the multimodal domain, being the first to analyze the domain-specific neurons in MLLMs. Compared with previous work on the interpretability of MLLM, such as those based on attention visualization Aflalo et al. (2022) or prompt-based probing Tao et al. (2024), our work stands out by providing some more fine-grained and solid quantitative analysis.
@@ 3 Method
In this section, we will introduce how to investigate the domain-specific neurons in MLLMs through domain activation probability entropy (DAPE). In Section 3.1, we define the activation of neurons in vision-language models. In Section 3.2, we introduce DAPE to reflect the specificity of neurons. Furthermore, to verify how post-projection embeddings are processed within the language model, we decode the hidden states layer by layer with logit lens, as discussed in Section 3.3.
@@ 3.2 Domain-Specific Neuron Selection
Our selection method is based on Tang et al. (2024). For each domain $D_{i},i=1,2,...,k$, we feed its image-text corpus into MLLM, and record the activated frequency of each neuron $u$ as well as the total token nums $N_{u,i}$. The activation probability of a neuron $u$ in domain $D_{i}$ is denoted as:
$p_{u,i}=\frac{M_{u,i}}{N_{u,i}},$ (4)
where $M_{u,i}$ refers to the activation frequency of neuron $u$ in domain $i$. We then denote the probability distribution of neuron $u$ across all domains as $P_{u}$:
@@ 3.3 Latent Embeddings Interpretation
As shown in Figure 4, the logit lens decodes the hidden states of the transformer’s intermediate layers into the distribution over the vocabulary, which can be used to interpret the model’s latent embeddings nostalgebraist (2020). Ideally, the decoded distribution converges monotonically toward the next token predicted by the model. And the results are so-called first-order or direct effect in some literature Gandelsman et al. (2023).
@@ 4.1.2 Dataset and Metrics
We select five datasets representing five different domains, namely, VQAv2 Goyal et al. (2017) for common scenes, PMC-VQA Zhang et al. (2023b) for Medical domain, DocVQA Mathew et al. (2021) for Document domain, LingoQA Marcu et al. (2023) for Auto Driving domain and RS-VQA Lobry et al. (2020) for Remote Sensing domain. For LingoQA, visual instruction for each question includes multiple images. More details can be found in Appendix C. We prepare image-question pairs of nearly the same token numbers for each domain during identifying, around 20 million tokens in LLaVA-NeXT. During evaluation, the scale of the validation set is aligned with LingoQA to make a fair comparison. For DocVQA, we report Average Normalized Levenshtein Similarity (ANLS) score Biten et al. (2019) followed by the official benchmark. For LingoQA, we use the score of Lingo-Judge Marcu et al. (2023) with the official implementation. For all other datasets, we report the top-1 accuracy (%) as the metric.
@@ 4.1.3 Implementation Details
We adhere to the default prompt templates from the official repository or the original paper during evaluation, with an additional role description for the auto-driving scenes. For more details, please refer to Appendix C. We perform the forward pass without padding or truncation during the identification process. When evaluating models across different datasets, we employ beam search with max_length of 512 and num_beams of 5 to generate answers. The temperature and length_penalty arguments are set as 0.9 and -1, respectively.
@@ 4.2.1 Distribution of Domain-specific Neurons
We identify domain-specific neurons using the method described in Section 3.2. Since neurons in different modules may have different activation patterns, as shown in Appendix D, we detected those domain-specific neurons module by module. Figure 5 shows the distribution of domain-specific neurons for each layer in each module of MLLMs.
Three-stage mechanism of LLM understanding multimodal features.
Two obvious turning points can be observed in both LLaVA-NeXT and InstructBLIP’s language model, one in the intermediate layer and the other near the output layer. Inspired by Zhao et al. (2024c), we thus propose a three-stage mechanism of LLM understanding multimodal features: 1) In the first several layers, projected features are further aligned with word space. Around the turning point, the multimodal features are embedded into a uniform representation space, where included domain-specific information needs to be processed by more domain-specific neurons. 2) Transitioning into the second phase, features are further generalized and understood by language models, where domain-specific neurons decrease sharply. 3)In the third stage, language models generate responses to the input, showing a rise of neurons specific to target tasks.
Our hypothesis aligns with the previous conclusion on smaller multimodal models like LiMBeR-BEIT Merullo et al. (2022), as Schwettmann et al. (2023) argue that outputs of the projection layer are further translated within the transformer after being merged with text embeddings. To further validate our hypothesis, we employ logit lens to visualize the transformation of multimodal features within language models in Section 4.2.3.
Domain-specific information in different semantic levels.
Domain-specific neurons are mainly distributed in shallow and intermediate layers within MLLMs’ vision encoders. Prior research discussed the correlation between the semantic level and layer depth, which found that more deep layers will focus on higher-level concepts in visual networks Xu et al. (2023); Raghu et al. (2021). In our settings, the document domain contains more low-level concepts, such as line and shape, while the remote sensing and medical domain may include more high-level concepts, like architectures and organs. Therefore, document neurons are mainly gathered in bottom layers close to the input end. Another interesting phenomenon is the rise of auto driving neurons near the output layer of InstructBLIP’s Q-Former, we conjecture this may reflect the struggle of model to understand the language instructions of auto driving domain.
Table 1 demonstrates the number of neurons in each domain. Remote sensing neurons have the largest proportion in LLaVA-NeXT’s vision encoder, MLP projector and language model, while in InstructBLIP, the domain owns most specific neurons are document, auto driving and auto driving separately. We argue that the number of specific neurons reflects the understanding ability of MLLM in the target domain, as more specific neurons may mean more demanding to process domain-specific information. In contrast, less specific neurons mean more generalized features in the target domain Tang et al. (2024). This also demonstrates a correlation between the training source and domain neuron distribution, as more data exposed during training resulting in less neurons specific for corresponding domain. In this way, we find that there exists a large visual gap between domains like remote sensing, document and medical, comparing the two domains left. Moreover, InstructBLIP seems less proficient in processing questions from auto driving, as neurons of this domain exhibit the highest number in Q-Former and LLM. There is also a similar pattern in its language model as for the auto driving domain. In other words, while visual features of auto driving domain can be processed well by existing vision encoder, the language instruction of this domain may be hard to handle for language model.
@@ 4.2.3 Case Study
To investigate how MLLM’s language model processes image tokens, we employ logit lens nostalgebraist (2020) to decode the hidden states of the language model’s intermediate layers into the probability of the next token across the vocabulary. As displayed in Figure 6, when feeding a remote sensing image-question pair into InstructBLIP, we get that the most likely token next to the second image token is ’</s>’, while the most likely token next of the last text token is the correct answer, ’no’. Interestingly, two place names, "Hermann" and "Baltimore", have appeared among the top token candidates when the image input is a remote sensing picture of New York. In multilingual literature, similar phenomena have also been observed. For instance, when Llama 2 receives the French token ’fleur’ in the input, the English concept ’__flower’ will appear in the intermediate distribution Wendler et al. (2024). This suggests that the decoded vocabulary distribution can to some extent reflect the semantic concepts understood by the language model. Despite this observation, we note that the decoded distribution of image tokens is far more sparse than text tokens; even in the output layer, the probability of the most likely next token ’</s>’ is lower than 40%. It indicates that projected tokens may be treated as a sparse mixture of concepts in the representation space instead of a simple word. We also demonstrate more cases of logit lens in different domains in Appendix E.
@@ 5 Conclusion
Acknowledgements
This work was supported by the National Key R&D Program of China (Grant No.2023YFF0725001); National Natural Science Foundation of China (Grant No.92370204); Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2023A03J0008), Education Bureau of Guangzhou Municipality; Guangdong Provincial Department of Education Project (Grant No.2024KQNCX028); Scientific Research Projects for the Higher-educational Institutions (Grant No.2024312096), Education Bureau of Guangzhou Municipality; Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.SL2024A03J01201), Education Bureau of Guangzhou Municipality; China Association for Science and Technology (Grant No.XMSB20240711064).
Appendix A Appendix
Appendix B Visual Domain Definition
@@
Appendix C Prompt Template
C.1 Instructions templates for VQA
@@
For instructions with options, we separate options in alphabetical order, as shown in Appendix C.2. $\star$ : A role description has been provided to help models better understand the tasks in auto driving. As shown below:
@@ C.2 Prompt Examples
We display the prompt format we use for evaluation in LLaVA-NeXT, as shown in Figure 8. The prompt for InstructBLIP come from direct format in Table C.1.
Appendix D Silent Neurons in MLLM’s Vision Encoder
@@
Appendix E Logit Lens Cases
@@
Appendix F Sensitivity and Scalability Analysis
To verify the robustness and scalability of our method, we further conducted the domain-specific neuron selection experiment at a different threshold of 5%, and complete analysis on llava-v1.6-vicuna-13b-hf. We report the results in Table 6, 7, 8 and 9.
Baseline Module VQAv2 PMC-VQA LingoQA DocVQA RS-VQA
LLaVA-NeXT Vision Encoder 1709 2103 1789 3046 2527
MLP Projector 106 120 128 87 102
LLM 9775 10839 11326 7445 12893
InstructBLIP Vision Encoder 1656 2868 2927 4402 5212
Q-Former 526 927 1925 902 949
LLM 4269 4303 8113 3350 8911
Table 6: The number of neurons in each domain in different modules of MLLMs at the threshold of 5%.
Baseline Module VQAv2 PMC-VQA LingoQA DocVQA RS-VQA
LLaVA-NeXT-13B Vision Encoder 65 224 156 430 438
MLP Projector 2 25 9 16 15
LLM 992 1276 2518 605 3289
Table 7: The number of neurons in each domain in different modules of llava-v1.6-vicuna-13b-hf at the threshold of 1%.
Model Deactivated Module(s) VQAv2 PMC-VQA LingoQA RS-VQA DocVQA
LLaVA-NeXT None 83.6 34.8 37.2 54.2 74.7
Vision Encoder 84.1 33.8 34.0 51.2 74.4
MLP Projector 84.1 34.8 35.9 54.2 74.7
LLM 83.6 33.8 36.2 48.3 75.4
All 84.1 31.8 33.5 50.2 75.0
Table 8: Accuracy (%) of LLaVA-NeXT-13B on selected domains with corresponding domain-specific neurons deactivated.
Token Type Layer 0 Layer 11 Layer 28 Layer 40
AEVP for Image Tokens 8.2734 8.3750 3.3359 3.6152
AEVP for Text Tokens 8.1797 8.4688 2.4375 2.2656
Table 9: Average Entropy of Vocab Probability (AEVP) of LLaVA-NeXT-13B in selected layers.